OpenStudy (anonymous):

Company ABC produces widgets. They have found that the cost, c (x), of making x widgets is a quadratic function in terms of x. The company also discovered that it costs $23 to produce 2 widgets, $55 to produce 4 widgets, and $247 to produce 10 widgets. What is the total cost of producing 8 widgets?

OpenStudy (anonymous):

i cantseem to get this in the right formula setup to solve? Can anyone help out?

OpenStudy (anonymous):

let C(x)=ax^2+bx+c then 23 =4a+2b+c ( when x=2) 55=16a+4b+c (when x=4) 247=100a+10b+c (whenx=10)

OpenStudy (anonymous):

by solving these equations for a,b,c the required function can be evaluated

OpenStudy (anonymous):

how do i solve with the 3 variables present?

OpenStudy (anonymous):

do i have to set up a matrix to solve this or what? i don't know where to go from here...help please